matt and meg comer are married and file a joint tax return. they do not have any children, matt works as a history professor at a local university and earns a salary of $65,850. meg works part time at the same university. she earns $31,750 a year. the couple does not itemized deductions. other than salary, the comer only other source of income is from the disposition of various capital assets (most stocks). (use the tax rate schedule, dividends and capital gains tax rates. ) (round final answer to the nearest whole dollar amount.)
a) what is the comer tax liability for 2018 if they report the following capital gains and losses for the year?
short term capital gains $9,150
short term capital losses $(2,150)
long term capital gains $15,290
long term capital losses $(6,290)
total tax liability

b) what is the comer tax liability for 2018 if they report the following capital gains and losses for the year?
short term capital gains $1,500
short term capital losses $0
long term capital gains $11,580
long term capital losses $(10,150)
total tax liability

can you please type it out
Answer #1

a) tax on income = 0.4(65850+31750) = $ 39040

On short term capital gains = 0.4(9150-2150) = 2800

For joint return the long term capital gains are not Taxable upto $77201

So tax liablity = 39040+ 2800 = $41840

b)tax on income same as a = 39040

short term capital gains 0.4(1500) = 600

Tax liability = $30640